question 2

the following data are available for a foundry operation started as a new company four years ago when the construction cost index was 125

current liabilities $170000

operating income $176200

NBV long term assets (end year 3) $687500

current assets $300000

GROSS BOOK VALUE $1100000

estimated total useful life 8 years

Age of assets 4 years

construction cost index end of year 4 150

of long term assets at historical cost

what is the ROI using current cost ?

1-12%

2-22.50%

3-(11.50)%

4-11.25%

5-16%
Answer #1

Calculations of Return on Investment

ROI = Net operating Income/ Total Value of Investments*100

Total Value of Investment = NBV of Long term Assets + Current Assets

Net Book Value of Long term Assets= Value at end of 3rd Year / Index at 4 Years ago * Index Value at end of 4th Year

= 687500/125*150= 825000

Total Value of Investment = 825000+300000

= 1125000

ROI = 176200/112500*100

=16%